# Conformal Prediction for Long-Tailed Classification

### v2

@@ Abstract
Many real-world classification problems, such as plant identification, have extremely long-tailed class distributions. In order for prediction sets to be useful in such settings, they should (i) provide good class-conditional coverage, ensuring that rare classes are not systematically omitted from the prediction sets, and (ii) be a reasonable size, allowing users to easily verify candidate labels. Unfortunately, existing conformal prediction methods, when applied to the long-tailed setting, force practitioners to make a binary choice between small sets with poor class-conditional coverage or sets with very good class-conditional coverage but that are extremely large. We propose methods with marginal coverage guarantees that smoothly trade off between set size and class-conditional coverage. First, we introduce a new conformal score function called prevalence-adjusted softmax that targets macro-coverage, a relaxed notion of class-conditional coverage. Second, we propose a new procedure that interpolates between marginal and class-conditional conformal prediction by linearly interpolating their conformal score thresholds. We demonstrate our methods on Pl@ntNet-300K and iNaturalist-2018, two long-tailed image datasets with 1,081 and 8,142 classes, respectively.
@@ 1 Introduction
Prediction sets are useful because they replace a single fallible point prediction with a set that is likely to contain the true label. In classification, prediction sets are most useful in settings with many classes, as they provide a selection ofcandidate labels that human decision makers can verify. Consider an amateur plant enthusiast who wants to identify a plant. The enthusiast struggles to identify plants on their own, but when presented with a short list of possible matches, it is easy for them to go through the list and select the correct species (e.g., by comparing their plant with images of potential matches).
Another key characteristic of plant identification is its extremely long-tailed class distribution. As shown in Figure˜1, there are thousands of images of common plants but only a handful for rare plants. Such skewed distributions also appear in animal identification and disease diagnosis. An added challenge is that we often care even more about identifying instances of the rare classes than the popular ones. In botany, scientists may want to prioritize the acquisition of examples of endangered plant species, which fall in the tail (Figure˜1). In medicine, catching the few cases of an aggressive cancer early matters more than perfectly classifying common benign lesions. In collaborative human-AI systems where a human generates labels based on AI recommendations and these labels are used to improve the predictive model in future training rounds, neglecting niche classes can accelerate “model collapse” (Shumailov et al., 2024), shrinking the model’s effective label space over time and degrading accuracy. This motivates pursuing prediction sets where all classes have a high probability of being correctly included in the prediction set. Beyond training good predictive models in this setting, an additional challenge for post-hoc uncertainty quantification is that most available examples of rare classes are used for model training, leaving these classes with few or zero holdout examples to use for calibrating uncertainty quantification methods.
@@
Formally, let $X\in\mathcal{X}$ be features with unknown label $Y\in{\mathcal{Y}}=\{1,2,\dots,|{\mathcal{Y}}|\}$. Our goal is to construct a set-generating procedure ${\mathcal{C}}:{\mathcal{X}}\to 2^{\mathcal{Y}}$, where $2^{\mathcal{Y}}$ denotes the set of all subsets of $\mathcal{Y}$}, with good class-conditional coverage. For $y\in{\mathcal{Y}}$, the class-conditional coverage of $\mathcal{C}$} for class $y$ is $\text{CondCov}({\mathcal{C}},y)={\mathbb{P}}(Y\in{\mathcal{C}}(X)\mid Y=y)$. This can be contrasted with marginal coverage, which is simply $\text{MarginalCov}({\mathcal{C}})={\mathbb{P}}(Y\in{\mathcal{C}}(X))$. To ensure that our prediction sets are useful for identifying instances of all classes, including rare ones, we aim for high class-conditional coverage across all classes. In addition to coverage, set size is also crucial, as large sets are often impractical. For instance, in plant identification, enthusiasts lack the time to review prediction sets containing hundreds of species. Ideally, we would generate prediction sets with both high class-conditional coverage and small size. Unfortunately, there is an inherent trade-off between small set sizes and class-conditional coverage in the long-tailed setting.
Conformal prediction (CP) provides methods guaranteed to achieve marginal or class-conditional coverage under no distributional assumptions. Standard CP, the most basic conformal prediction method, yields small sets but only guarantees marginal coverage and often has poor class-conditional coverage for some classes. However, approaches targeting class-conditional coverage struggle in scenarios where many classes have few examples. In such settings, Classwise CP (a Mondrian conformal variant Vovk et al., 2005) and rank-calibrated class-conditional CP (Shi et al., 2024) produce large sets, and Clustered CP (Ding et al., 2023) effectively defaults to Standard CP for rare classes that it is unable to assign to a cluster.
Our goal is to create prediction sets that optimize class-conditional coverage, without sacrificing reasonable set sizes or marginal coverage. We approach this in two ways.
@@
This emphasizes coverage of more frequent classes. We derive the form of the prediction set that optimally trades off set size and macro-coverage under oracle knowledge of the underlying distribution and design a conformal score function, called prevalence-adjusted softmax ($\mathsf{PAS}$), which approximates these oracle optimal sets given an imperfect classifier $\hat{p}(y|x)$ and estimated label distribution $\hat{p}(y)$.
@@ 1.1 Related work
Conformal prediction provides a way to construct prediction sets with coverage guarantees given a calibration dataset that is exchangeable with the test point (Vovk et al., 2005; Angelopoulos and Bates, 2023). We are specifically interested in class-conditional coverage, which is difficult to achieve in a useful way when many classes have very few calibration examples, as is the case in long-tailed settings. Previous works on class-conditional conformal focus on easy settings with at most ten classes (Shi et al., 2013; Löfström et al., 2015; Hechtlinger et al., 2018; Sadinle et al., 2019), or the harder setting of many classes with some class imbalance but still lacking a truly long tail (Ding et al., 2023; Shi et al., 2024).
@@
Prediction sets should achieve the desired coverage while being as small as possible, so as to be maximally informative. In general, the set-generating procedures that optimally navigate the coverage-size trade-off depend on the density of the underlying distribution (Lei et al., 2013; Lei and Wasserman, 2014; Vovk et al., 2016; Sadinle et al., 2019). Although the true density is unknown in practice, these theoretically optimal sets serve as guidelines for designing effective conformal score functions or new conformal procedures for various target quantities, such as $X$-conditional coverage (e.g., APS from Romano et al., 2020, RAPS from Angelopoulos et al., 2021, SAPS from Huang et al., 2023, CQC from Cauchois et al., 2021) or size (Denis and Hebiri, 2017; Kiyani et al., 2024). In Approach I, we construct a score function inspired by the oracle sets that optimally trade-off set size and macro-coverage, a coverage target that has not been previously explored in this context.
@@
Many real-world classification problems exhibit long-tailed distributions, a challenge particularly pronounced in fine-grained visual categorization, a field that has recently attracted substantial interest, especially in biodiversity (Van Horn et al., 2015, 2018; Garcin et al., 2021; Wang et al., 2022). The goal is to classify images into highly specific subcategories (e.g., plant species), which often differ only subtly. While extensive research has focused on improving class-conditional top-$1$ or top-$k$ accuracy for such tasks (Russakovsky et al., 2015; Lapin et al., 2015; Liu et al., 2019; Garcin et al., 2021; Zhang et al., 2023), less attention has been devoted to constructing high-quality prediction sets in long-tailed settings, beyond the naive approach of selecting the top-$k$ predictions. While methods like logit adjustment for long-tail learning (Menon et al., 2021) and focal loss for dense object detection (Lin et al., 2017) address class imbalance through reweighting or output adjustment to improve classification performance, designing robust prediction sets under long-tailed distributions remains understudied. Addressing this gap is critical for systems like Pl@ntNet (Joly et al., 2014), a plant identification app where users upload images and receive candidate species matches.
@@ 1.2 Preliminaries
For a positive integer $n$, let $n]:=\{1,\ldots,n\}$. Let ${\mathcal{D}}_{\mathrm{cal}}=\left\{(X_{i},Y_{i})\text{ for }i\in[n]\right\}$ be a calibration set , where $(X_{i},Y_{i})$ for $i=1,\dots,n$ are exchangeable with the test point $(X_{n+1},Y_{n+1})$, where the label $Y_{n+1}$ is unknown. For a class $y\in{\mathcal{Y}}$, we use ${\mathcal{I}}_{y}=\{i\in[n]:Y_{i}=y\}$ to denote the set of calibration points with label $y$ and $n_{y}=\left|{\mathcal{I}}_{y}\right|$ the number of calibrations points with label $y$. Let $\alpha\in[0,1]$ be a user-specified probability of miscoverage. We use $s:{\mathcal{X}}\times{\mathcal{Y}}\to\mathbb{R}$ to denote a conformal score function, where smaller values of $s(x,y)$ indicate that the pair $(x,y)$ conforms better with previously seen data.
@@
We focus on the split conformal setting in which the score function $s$ is constructed using a predictive model trained on data separate from the calibration set ${\mathcal{D}}_{\mathrm{cal}}$. We use the softmax conformal score function, defined as $s_{\mathsf{softmax}}(x,y)=1-\hat{p}(y|x)$, where $\hat{p}(y|x)$ is the predicted probability for class $y$ for input $x$ obtained from the softmax output of a trained neural network. Our methods can be readily adapted to other scoring functions but we focus on this score because alternatives like APS (Romano et al., 2020) and its variants produce larger sets and are primarily designed for $X$-conditional coverage, which is outside our scope.
@@
Classwise conformal prediction constructs sets as ${\mathcal{C}}_{\textsc{Classwise}}(X):={\mathcal{C}}(X;\hat{\mathbf{q}}_{\textsc{Classwise}}(s,{\mathcal{D}}_{\mathrm{cal}}))$ where the $y$-th entry of $\hat{\mathbf{q}}_{\textsc{Classwise}}(s,{\mathcal{D}}_{\mathrm{cal}})$ is
@@
We explicitly describe the meta-algorithm for conformal prediction in Algorithm˜1. The existing methods, Standard and Classwise, and the methods we will propose in the next section are instantiations of this meta-algorithm for the score functions and threshold functions described in Table˜1.
@@ 2 Methods
We take two approaches to the problem of simultaneously achieving reasonable set sizes and reasonable class-conditional coverage in the long-tailed classification setting, each leading to a method. Approach I targets the weaker guarantee of macro-coverage. From this, we derive an oracle set that optimally balances set size and macro-coverage. We then define a conformal score function ($\mathsf{PAS}$) by replacing the true conditional density with its estimate. We also consider an extension ($\mathsf{WPAS}$) that prioritizes coverage of user-specified classes. Approach II addresses the trade-off by interpolating between class-conditional and marginal score quantiles, leading to a simple new procedure (Interp-Q). We defer the formal statements of propositions and their proofs to Appendix B.
@@ 2.1 Approach I: Targeting (weighted) macro-coverage via a new score
$\min_{{\mathcal{C}}:\mathcal{X}\mapsto 2^{\mathcal{Y}}}{\mathbb{E}}[|{\mathcal{C}}(X)|]\quad\text{subject to }\text{MacroCov}({\mathcal{C}})\geq 1-\alpha,$ (8)
@@
where $\text{MacroCov}({\mathcal{C}})$ is defined in (1).
@@
for some threshold $t$ that depends on $\alpha$ or $\kappa$, respectively.
The key takeaway from this proposition is that thresholding on $p(y|x)/p(y)$ optimally balances macro-coverage and expected set size. Specifically, among set-generating procedures with a given expected set size, none achieve better macro-coverage than thresholding on $p(y|x)/p(y)$. Similarly, for a fixed macro-coverage, no other procedure yields a smaller expected set size. We contrast this with the solution to the more classical problem of minimizing the expected set size subject to marginal or class-conditional coverage, which is given by thresholding on $p(y|x)$ (Neyman and Pearson, 1933; Sadinle et al., 2019).
@@
and $\mathsf{PAS}$ stands for prevalence-adjusted softmax. By setting t-t to be the Standard CP $\hat{q}$ from (4) using $s_{\mathsf{PAS}}$, $\smash{\widehat{{\mathcal{C}}}}$ inherits the marginal coverage guarantee of Standard. In summary, the first method we propose is simply running Standard CP with the $\mathsf{PAS}$ score function (which we will refer to as Standard with $\mathsf{PAS}$), as this achieves the desired marginal coverage guarantee while (approximately) optimally trading off set size and macro-coverage.
@@
Recall that macro-coverage is the unweighted average of the class-conditional coverages, and the $\mathsf{PAS}$ score function is designed to optimize macro-coverage among all set-generating procedures with a certain expected set size. However, in some settings we may wish to instead optimize for a weighted average of the class-conditional coverages (e.g., because it is more important to cover some classes than others). Given user-chosen class weights $\omega(y)$ for $y\in{\mathcal{Y}}$ that sum to one, we can similarly define the $\omega$-weighted macro-coverage as
@@
for some threshold $t$ that depends on $\omega$ and $\alpha$ or $\kappa$, respectively.

### v3

# arXiv:2507.06867v3[stat.ML] 26 Feb 2026

@@ Abstract
Many real-world classification problems, such as plant identification, have extremely long-tailed class distributions. In order for prediction sets to be useful in such settings, they should (i) provide good class-conditional coverage, ensuring that rare classes are not systematically omitted from the prediction sets, and (ii) be a reasonable size, allowing users to easily verify candidate labels. Unfortunately, existing conformal prediction methods, when applied to the long-tailed setting, force practitioners to make a binary choice between small sets with poor class-conditional coverage or sets that have very good class-conditional coverage but are extremely large. We propose methods with marginal coverage guarantees that smoothly trade off set size and class-conditional coverage. First, we introduce a new conformal score function called prevalence-adjusted softmax that optimizes for macro-coverage, defined as the average class-conditional coverage across classes. Second, we propose a new procedure that interpolates between marginal and class-conditional conformal prediction by linearly interpolating their conformal score thresholds. We demonstrate our methods on Pl@ntNet-300K and iNaturalist-2018, two long-tailed image datasets with 1,081 and 8,142 classes, respectively.
@@ 1 Introduction
Prediction sets are useful because they replace a single fallible point prediction with a set that is likely to contain the true label. In classification, prediction sets are most useful in settings with many classes, as they narrow down the label space to a set of labels that human decision makers can then verify. Consider an amateur plant enthusiast who wants to identify a plant. The enthusiast struggles to identify plants on their own, but when presented with a short list of possible matches, it is easy for them to go through the list and select the correct species (e.g., by comparing their plant with images of potential matches).
Another key characteristic of plant identification is its extremely long-tailed class distribution. As shown in Figure 1, there are thousands of images of common plants but only a handful for rare plants. Such skewed distributions also appear in animal identification and disease diagnosis. An added challenge is that we often care even more about identifying instances of the rare classes than the popular ones. In botany, scientists may want to prioritize the acquisition of examples of endangered plant species, which fall in the tail (Figure 1). In medicine, catching the few cases of an aggressive cancer early matters more than perfectly classifying common benign lesions. In collaborative human-AI systems where a human generates labels based on AI recommendations and these labels are used to improve the predictive model in future training rounds, neglecting niche classes can accelerate “model collapse” (Shumailov et al., 2024), shrinking the model’s effective label space over time and degrading accuracy. This motivates pursuing prediction sets where all classes have a high probability of being correctly included in the prediction set. Beyond training good predictive models in this setting, an additional challenge for post-hoc uncertainty quantification is that most available examples of rare classes are used for model training, leaving these classes with few or zero holdout examples to use for calibrating uncertainty quantification methods.
@@
Formally, let $X\in\mathcal{X}$ be features with unknown label $Y\in{\mathcal{Y}}=\{1,2,\dots,|{\mathcal{Y}}|\}$. Our goal is to construct a set-generating procedure ${\mathcal{C}}:{\mathcal{X}}\to 2^{\mathcal{Y}}$, where $2^{\mathcal{Y}}$ denotes the set of all subsets of $\mathcal{Y}$}, with good class-conditional coverage. For $y\in{\mathcal{Y}}$, the class-conditional coverage of $\mathcal{C}$} for class $y$ is $\text{CondCov}({\mathcal{C}},y)={\mathbb{P}}(Y\in{\mathcal{C}}(X)\mid Y=y)$. This can be contrasted with marginal coverage, which is simply $\text{MarginalCov}({\mathcal{C}})={\mathbb{P}}(Y\in{\mathcal{C}}(X))$. To ensure that our prediction sets are useful for identifying instances of all classes, including rare ones, we aim for high class-conditional coverage across all classes. In addition to coverage, set size is also crucial, as large sets are often impractical. For instance, in plant identification, users lack the time to review prediction sets containing hundreds of species. Ideally, we would generate prediction sets with both high class-conditional coverage and small size. Unfortunately, there is an inherent trade-off between small set sizes and class-conditional coverage in the long-tailed setting.
Conformal prediction (CP) provides methods guaranteed to achieve marginal or class-conditional coverage under no distributional assumptions. Standard CP, the most basic conformal prediction method, yields small sets but only guarantees marginal coverage and often has poor class-conditional coverage for some classes. However, approaches targeting class-conditional coverage struggle in scenarios where many classes have few examples. In such settings, Classwise CP (an instantiation of Mondrian conformal; see Vovk et al., 2005) and rank-calibrated class-conditional CP (Shi et al., 2024) produce large sets, and Clustered CP (Ding et al., 2023) effectively defaults to Standard CP for rare classes that it is unable to assign to a cluster.
Objective. Our goal is to produce prediction sets that maintain marginal coverage while striking a more useful trade-off between set size and class conditional coverage compared to Standard or Classwise CP. We approach this in two ways.
@@
This (over)emphasizes coverage of more frequent classes. We derive the form of the prediction set that optimally trades off set size and macro-coverage under oracle knowledge of the underlying distribution and design a conformal score function, called prevalence-adjusted softmax ($\mathsf{PAS}$), which approximates these oracle optimal sets given an imperfect classifier $\hat{p}(y|x)$ and estimated label distribution $\hat{p}(y)$.
@@ 1.1 Related work
Conformal prediction provides a way to construct prediction sets with coverage guarantees given a calibration dataset that is exchangeable with the test point (Vovk et al., 2005; Angelopoulos & Bates, 2023). We are specifically interested in class-conditional coverage, which is difficult to achieve in a useful way when many classes have very few calibration examples, as is the case in long-tailed settings. Previous works on class-conditional conformal focus on easy settings with at most ten classes (Shi et al., 2013; Löfström et al., 2015; Hechtlinger et al., 2018; Sadinle et al., 2019), or the harder setting of many classes with some class imbalance but still lacking a truly long tail (Ding et al., 2023; Shi et al., 2024).
@@
Prediction sets should achieve the desired coverage while being as small as possible, so as to be maximally informative. In general, the set-generating procedures that optimally navigate the coverage–size trade-off depend on the density of the underlying distribution (Lei et al., 2013; Lei & Wasserman, 2014; Vovk et al., 2016; Sadinle et al., 2019). Although the true density is unknown in practice, these theoretically optimal sets serve as guidelines for designing effective conformal score functions or new conformal procedures for various target quantities, such as $X$-conditional coverage (e.g., APS from Romano et al., 2020, RAPS from Angelopoulos et al., 2021, SAPS from Huang et al., 2023, CQC from Cauchois et al., 2021) or size (Denis & Hebiri, 2017; Kiyani et al., 2024). In Approach I, we construct a score function inspired by the oracle sets that optimally trade-off set size and macro-coverage, a coverage target that has not been previously explored in this context.
@@
Many real-world classification problems exhibit long-tailed distributions, a challenge particularly pronounced in fine-grained visual categorization, a field that has recently attracted substantial interest, especially in biodiversity (Van Horn et al., 2015; 2018; Garcin et al., 2021; Wang et al., 2022). The goal is to classify images into highly specific subcategories (e.g., plant species), which often differ only subtly. While extensive research has focused on improving class-conditional top-$1$ or top-$k$ accuracy for such tasks (Russakovsky et al., 2015; Lapin et al., 2015; Liu et al., 2019; Garcin et al., 2021; Zhang et al., 2023), less attention has been devoted to constructing high-quality prediction sets in long-tailed settings, beyond the naive approach of selecting the top-$k$ predictions. While methods like logit adjustment for long-tail learning (Menon et al., 2021) and focal loss for dense object detection (Lin et al., 2017) address class imbalance through reweighting or output adjustment to improve classification performance, designing robust prediction sets under long-tailed distributions remains understudied. Addressing this gap is critical for systems like Pl@ntNet (Joly et al., 2014), a plant identification app where users upload images and receive candidate species matches.
@@ 1.2 Preliminaries
For a positive integer $n$, let $n]:=\{1,\ldots,n\}$. Let ${\mathcal{D}}_{\mathrm{cal}}=\left\{(X_{i},Y_{i})\text{ for }i\in[n]\right\}$ be a calibration set, where $(X_{i},Y_{i})$ for $i=1,\dots,n$ are exchangeable with the test point $(X_{n+1},Y_{n+1})$, where the label $Y_{n+1}$ is unknown. For a class $y\in{\mathcal{Y}}$, we use ${\mathcal{I}}_{y}=\{i\in[n]:Y_{i}=y\}$ to denote the set of calibration points with label $y$ and $n_{y}=\left|{\mathcal{I}}_{y}\right|$ the number of calibration points with label $y$. Let $\alpha\in[0,1]$ be a user-specified probability of miscoverage. We use $s:{\mathcal{X}}\times{\mathcal{Y}}\to\mathbb{R}$ to denote a conformal score function, where smaller values of $s(x,y)$ indicate that the pair $(x,y)$ conforms better with previously seen data.
@@
We focus on the split conformal setting in which the score function $s$ is constructed using a predictive model trained on data separate from the calibration set ${\mathcal{D}}_{\mathrm{cal}}$. We use the softmax conformal score function, defined as $s_{\mathsf{softmax}}(x,y)=1-\hat{p}(y|x)$, where $\hat{p}(y|x)$ is the predicted probability for class $y$ for input $x$ obtained from the softmax output of a trained neural network. This is also known as the Least Ambiguous Classifier (LAC) score (Sadinle et al., 2019). Our methods can be readily adapted to other scoring functions but we focus on this score because alternatives like APS (Romano et al., 2020) and its variants produce larger sets and are primarily designed for $X$-conditional coverage, which is outside our scope.
@@
Classwise conformal prediction constructs sets as ${\mathcal{C}}_{\textsc{Classwise}}(X):={\mathcal{C}}(X;\hat{\mathbf{q}}_{\textsc{Classwise}})$ where the $y$-th entry of $\hat{\mathbf{q}}_{\textsc{Classwise}}$ is
@@
We explicitly describe the meta-algorithm for conformal prediction in Algorithm 1. The existing methods, Standard and Classwise, and the methods we will propose in the next section are instantiations of this meta-algorithm for the score functions and threshold functions described in Table 1.
@@ 2 Methods
We take two approaches to the problem of simultaneously achieving reasonable set sizes and reasonable class-conditional coverage in the long-tailed classification setting, each leading to a method. Approach I targets the weaker objective of macro-coverage. From this, we derive an oracle set that optimally balances set size and macro-coverage. We then define a conformal score function ($\mathsf{PAS}$) by replacing the true conditional density with its estimate. We also consider an extension ($\mathsf{WPAS}$) that prioritizes coverage of user-specified classes. Approach II addresses the trade-off by interpolating between class-conditional and marginal score quantiles, leading to a simple new procedure (Interp-Q). We defer the formal statements of propositions and their proofs to Appendix B.
@@ 2.1 Approach I: Targeting (weighted) macro-coverage via a new score
$\min_{{\mathcal{C}}:\mathcal{X}\mapsto 2^{\mathcal{Y}}}{\mathbb{E}}[|{\mathcal{C}}(X)|]\quad\text{subject to }\text{MacroCov}({\mathcal{C}})\geq\beta,$ (8)
@@
where $\text{MacroCov}({\mathcal{C}})$ is defined in (1) and $\beta\geq 0$ and $\kappa\geq 0$.
@@
for some threshold $t$ that depends on $\beta$ or $\kappa$, respectively.
The key takeaway from this proposition is that thresholding on $p(y|x)/p(y)$ optimally balances macro-coverage and expected set size. Specifically, among set-generating procedures with a given expected set size, none achieve better macro-coverage than thresholding on $p(y|x)/p(y)$. Similarly, for a fixed macro-coverage, no other procedure yields a smaller expected set size. We contrast this with the solution to the more classical problem of minimizing expected set size subject to marginal or class-conditional coverage, which is given by thresholding on $p(y|x)$ (Neyman & Pearson, 1933; Sadinle et al., 2019).
@@
and $\mathsf{PAS}$ stands for prevalence-adjusted softmax. By setting t-t as the Standard CP $\hat{q}$ from (4) using $s_{\mathsf{PAS}}$ as the score, $\smash{\widehat{{\mathcal{C}}}}$ inherits the marginal coverage guarantee of Standard CP. In summary, the first method we propose is simply running Standard CP with the $\mathsf{PAS}$ score function (which we will refer to as Standard with $\mathsf{PAS}$), as this achieves the desired marginal coverage guarantee while (approximately) optimally trading off set size and macro-coverage. We emphasize that $\mathsf{PAS}$ aims to better handle this trade-off but does not directly target a macro-coverage guarantee.
@@
Recall that macro-coverage is the unweighted average of the class-conditional coverages, and the $\mathsf{PAS}$ score function is designed to optimize macro-coverage among all set-generating procedures with a certain expected set size. However, in some settings, we may instead wish to optimize for a weighted average of the class-conditional coverages (e.g., because it is more important to cover some classes than others). Given user-chosen class weights $\omega(y)$ for $y\in{\mathcal{Y}}$ that sum to one, we can similarly define the $\omega$-weighted macro-coverage as
@@
for some threshold $t$ that depends on $\omega$ and $\beta$ or $\kappa$, respectively.